OpenStudy (anonymous):

how to solve the equation 5+6y-8=5y-3-y ?

OpenStudy (anonymous):

2y=0 y=0

OpenStudy (mathstudent55):

5 + 6y - 8 = 5y - 3 - y First combine like terms on each side: 6y - 3 = 4y - 3 Now subtract 4y from both sides: 2y - 3 = -3 Add 3 to both sides: 2y = 0 Divide both sides by 2: y = 0

OpenStudy (koikkara):